Compute the probability of event E if the odds in favor of E are 3 to 29.
Answer #1

please rate my answer,if it is helpful,thankyou

The odd of an event is the ratio of favorable outcomes to unfavorable outcomes.

total outcomes =3+29=32

favour outcomes =3

P(E)=3/32